OpenStudy (anonymous):

Simplify: ((1/x)-(1/5))/((1/x^2)-(1/25)) I've been stumped on this for a while.

OpenStudy (anonymous):

same denominator. top becomes: (5-x)/(5x) bottom becomes: (25-x^2)/25x^2 (a/b)/(c/d)=ad/bc so you get ((25x^2)(5-x))/(5x(25-x^2) can u simplify that

OpenStudy (anonymous):

you get (5x(5-x))/(25-x^2) (a^2-b^2)=(a-b)(a+b) so you get 5x/(5+x) final answer

OpenStudy (radar):